### v3

Abstract
Language model post-training is applied to refine behaviors and unlock new skills across a wide range of language models, but open recipes for applying these techniques lag behind proprietary ones. The underlying training data and recipes for post-training are simultaneously the most important pieces of the puzzle and the portion with the least transparency. To bridge this gap, we introduce Tülu 3, a family of fully-open state-of-the-art post-trained models, alongside its data, code, and training recipes, serving as a comprehensive guide for modern post-training techniques. Tülu 3, which builds on Llama 3.1 base models, achieves results surpassing the instruct versions of Llama 3.1, Qwen 2.5, Mistral, and even closed models such as GPT-4o-mini and Claude 3.5-Haiku. The training algorithms for our models include supervised finetuning (SFT), Direct Preference Optimization (DPO), and a novel method we call Reinforcement Learning with Verifiable Rewards (RLVR). With Tülu 3, we build a multi-task evaluation scheme for post-training with development and unseen evaluations, standard benchmark implementations, and substantial decontamination of existing open datasets on said benchmarks. We conclude with analysis and discussion of training methods that did not reliably improve performance.
The Tülu 3 release includes model weights, a demo, and the complete recipe — datasets for diverse core skills, a robust toolkit for data curation and evaluation, the training code and infrastructure, and, most importantly, a detailed report for reproducing and further adapting the Tülu 3 approach to more domains.
Tulu 3 8B:]Llama-3.1-Tulu-3-8B
[…]
Tulu 3 70B:]Llama-3.1-Tulu-3-70B
[…]
Tulu 3 405B:]Llama-3.1-Tulu-3-405B
[…]
Tulu 3 Data:]tulu-3-datasets-673b8df14442393f7213f372
[…]
Tulu 3 Code:]open-instruct
[…]
Tülu 3 Eval:]olmes
[…]
Demo:]playground.allenai.org
Contents
| Model Checkpoints | ||
| Stage | Llama 3.1 8B | Llama 3.1 70B |
| Base Model | meta-llama/Llama-3.1-8B | meta-llama/Llama-3.1-70B |
| SFT | allenai/Llama-3.1-Tulu-3-8B-SFT | allenai/Llama-3.1-Tulu-3-70B-SFT |
| DPO | allenai/Llama-3.1-Tulu-3-8B-DPO | allenai/Llama-3.1-Tulu-3-70B-DPO |
| Final Models | ||
| (RLVR) | allenai/Llama-3.1-Tulu-3-8B | |
| RM: allenai/Llama-3.1-Tulu-3-8B-RM | allenai/Llama-3.1-Tulu-3-70B | |
| Stage | Llama 3.1 405B | |
| Base Model | meta-llama/Llama-3.1-405B | |
| SFT | allenai/Llama-3.1-Tulu-3-405B-SFT | |
| Final Model | ||
| (DPO) | allenai/Llama-3.1-Tulu-3-405B |
| Codebases / Tools | |
| Type |
| Training | allenai/open-instruct |
| Tülu 3 Eval | allenai/olmes |
| Decontamination | allenai/open-instruct/tree/main/decontamination |
| Preference Data Inference | allenai/birr |
| Instruction Datasets | ||
| Type | Domain |
| Full mix | General | allenai/tulu-3-sft-mixture |
| Task Specific | Precise Instruction Following | allenai/tulu-3-sft-personas-instruction-following |
| Subsets | MATH | allenai/tulu-3-sft-personas-math |
| Grade School Math | allenai/tulu-3-sft-personas-math-grade | |
| Python Code | allenai/tulu-3-sft-personas-code | |
| Preference Mixes | |
| Model |
| Llama 3.1 70B | allenai/llama-3.1-tulu-3-70b-preference-mixture |
| Llama 3.1 8B | allenai/llama-3.1-tulu-3-8b-preference-mixture |
| Specific Preference Datasets | |
| Domain |
| Precise Instruction Following | allenai/tulu-3-pref-personas-instruction-following |
| General | allenai/tulu-3-sft-prompts-ultrafeedback |
| General | allenai/tulu-3-wildchat-ultrafeedback |
| RL with Verifiable Rewards Training Datasets | |
| Domain |
| Full Mix | allenai/RLVR-GSM-MATH-IF-Mixed-Constraints |
| GSM Only | allenai/RLVR-GSM |
| MATH Only | allenai/RLVR-MATH |
| IFeval Only | allenai/RLVR-IFeval |
1 Introduction
“Just as the camel shares its burdens with others in the caravan, the wise share their insights to lighten the load of ignorance.” – Proverb generated by Tülu 3.
Post-training — the collection of techniques including instruction tuning, reinforcement learning from human feedback, and other types of finetuning — has become a crucial step in building frontier language models (OpenAI2024; Anthropic2024), yet developments to these techniques are frequently not accompanied by open resources and recipes. Fully open source counterparts (e.g., Tülu 2 (ivison2023camels) and Zephyr- (tunstall2023zephyr)) often rely on simpler-to-implement and cheaper pipelines and have become outdated on many metrics.
To close the gap between open and closed post training, we introduce Tülu 111A tülu is a hybrid camel bred between Bactrian camel and dromedary: https://en.wikipedia.org/wiki/Hybrid_camel. 3, a family of open state-of-the-art post-trained models, alongside all of the data, training recipes, code, infrastructure, and evaluation framework. Integrating partial details from proprietary methods with novel techniques and established academic research, Tülu 3 pushes the boundaries of research in post-training. The advancements of Tülu 3 are attributed to Tülu 3 Data, new permissively licensed training datasets targeting core skills, Tülu 3 Eval, an evaluation suite and tools to establish clear performance goals and guide improvement through training stages, and Tülu 3 Recipe, an advanced multi-stage training pipeline incorporating new algorithmic advancements in reinforcement learning, cutting-edge infrastructure, and rigorous experimentation to optimize data mixes, methods, and parameters across various training stages.
In order to build Tülu 3, we identify a set of core skills to improve after training (e.g., reasoning, math, coding, safety, precise instruction following, knowledge recall, etc.) and build an evaluation framework to establish clear performance goals and guide model improvement over a selection of development and unseen tasks. Tülu 3 benefits significantly from leveraging publicly available open data, generating diverse, skill-specific synthetic data at various training stages, and aggressively decontaminating them against our evaluation suite.
The Tülu 3 training recipe involves multiple stages, with each stage building upon the previous model and focusing on different types of data — namely, prompt-completion instances for supervised finetuning, preferences for preference tuning, or verifiable rewards for reinforcement learning. Our methodology facilitates identifying skill deficiencies and refining the data mix, methods and parameters, ensuring a balanced performance of core skills across the training process. Through rigorous, principled experimentation, we determine the best data mix for supervised finetuning, resulting in the Tülu 3 SFT checkpoint. Leveraging recent advances in preference tuning, we then train a model over carefully curated on-policy preference data from comparing Tülu 3 SFT completions against outputs from other language models. Furthermore, we introduce a new final finetuning stage – Reinforcement Learning with Verifiable Rewards (RLVR) - which employs a novel RL objective tailored to enhance specific skills with verifiable answers, such as mathematics and precise instruction following.
Our best performing recipe yields Tülu 3 models that outperform the state-of-the-art post-trained open-weight models of the same size such as Llama 3.1 Instruct (dubey2024llama) or Mistral-Instruct (mistral2024ministraux), and at the large 70B size Tülu matches the offerings of closed providers such as Claude 3.5 Haiku and GPT-4o mini. Furthermore, at 405B size our model performs competitively against DeepSeek v3 (deepseekai2024deepseekv3technicalreport) and GPT 4o (11-24).
In summary, Tülu 3 represents a family of state-of-the-art open language models, featuring a modern post-training framework with fully open-source data Tülu 3 Data, evaluation Tülu 3 Eval, training code Tülu 3 Code and development recipes Tülu 3 Recipe. Here are a few key contributions from the development of Tülu:
Extensive guidance and tooling for evaluation, decontamination, and recipe design,
Scaled, new synthetic instruction datasets,
Scaling preference data with on-policy generations,
Reinforcement learning with verifiable rewards, an RL-based method that only gets a reward if the model’s completions are verified to be correct, and
Advanced infrastructure, details, and code to facilitate the successful implementation of large models.